OpenStudy (anonymous):

the sum of measures of the interior angles of a regular polygon is 540 degrees how many sides does the polygon have

OpenStudy (anonymous):

the sum of angles of ANY Convex polygon = 180*(Number of Vertexes -2)

OpenStudy (anonymous):

Also the numb. of vertexes is equal number the of sides

OpenStudy (anonymous):

180*n = 540 What is n ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

Then the number of sides is really n+2 = 5